I need help figuring this out.
The area of a circle 28 square meters. What is its cicumference? (Use A=nr^2 and C=2nr to figure out the problem)

Answer by rfer(16322)

A=Pir^2
28=3.14(r)^2
(r)^2=28/3.14
r^2=8.9
sq rt of r=2.97
-----------------
C=2Pir
C=2(3.14)(2.97)
C=18.75